How many cu m of butterfat must be removed from 500 cu m of milk containing 10% butterfat to reduce its content to 4% butterfat?
Answer by jainenderkapoor(61)

It is a very good question
Let us start with it.
How much milk we are having -- ? 500 cu m
And the percentage of butterfat is 10%
i:e butterfat = (10/100)*500 = 50
Now we want to remove some amount of butterfat so that the percentage of butterfat in the remaining liquid is 4%
NOW if we remove butterfat = x cu m then the same quantity will be removed from the milk -- IS IT CLEAR
So we have remaining quantity of buttefat = (50 - x)
and the remaining quantity of milk = (500 - x)
Is it clear
Now we can write the equation --
(50 - x) = 4% of (500 - x)
(50 - x) = 4/100 * (500 - x)
(50 - x) = 1/25 * (500 - x)
25*(50 - x) = (500 - x)
25*50 - 25x = 500 - x
1250 - 25x = 500 -x
1250 - 500 = -x + 25x
750 = 24 x
24 x = 750
x = 750/24 = 31.25
SO 31.25 cu m of butterfat is to be removed so that the content is 4% of the remaining liquid.
HOPE it is clear to you. In case you have any other doubt you can contact me on kapoorjai1@rediffmail.com
I am available for Online Tutoring Also.